OpenStudy (anonymous):

2. If f(x)=2x^2-5x, find the values of k if f^-1(3)=k @iambatman @wio

OpenStudy (anonymous):

Start off by finding the inverse of f(x)

OpenStudy (anonymous):

Remember that: \[ f^{-1}(f(x)) = x \]

OpenStudy (anonymous):

So when \(x=k\), we can say \(f(x)=3\).

OpenStudy (anonymous):

Which also means that \(f(k)=3\)

OpenStudy (anonymous):

Yeah, that's good

OpenStudy (anonymous):

I tried by i got \[f^-1(x)=\sqrt{2x+10}\]

OpenStudy (anonymous):

but*

OpenStudy (anonymous):

\[f(k)=2k^2-5k=3\]

OpenStudy (anonymous):

So you need to solve for \(k\) here. It's a quadratic equation.

OpenStudy (anonymous):

\[ 2k^2-5k-3=0 \]

OpenStudy (anonymous):

k=5 and k=3/2

OpenStudy (anonymous):

Do you know how to solve quadratic equations

OpenStudy (anonymous):

Your inverse function is wrong, \[f(x)=2x^2-5x\] The inverse of this function is \[f^{-1} (x)= \frac{ 1 }{ 4 } (5 \pm \sqrt{8x+25})\] \[f^{-1} (3) = \frac{ 1 }{ 4 }(5\pm \sqrt{8(3)+25})\] \[f^{-1} (3) = \frac{ 1 }{ 4 }(5 \pm 7) = k\]

OpenStudy (anonymous):

Here is how you would find and inverse: \[ f(x) =ax^2+bx+c\\ 0=ax^2+bx+c-f(x) \]Then you would use quadratic equation, with \(c-f(x)\) substituted in for what would be \(c\).\[ x = \frac{-b\pm\sqrt{b^2-4a(c-f(x))}}{2a} \]By definition \(f^{-1}(f(x))=x\), and we clearly have an expression of \(f(x)\) that equals \(x\), the only problem is that the \(\pm\) means it is not a function. We have two options:\[ f^{-1}(x)=\frac{-b\color{red}+\sqrt{b^2-4a(c-x)}}{2a}\\ f^{-1}(x)=\frac{-b\color{red}-\sqrt{b^2-4a(c-x)}}{2a} \]

OpenStudy (anonymous):

thnx a lot @iambatman @wio